A company purchased $3000 of merchandise on july 5 with terms 2/10, n/30. On july 7, it returned $330 worth of merchandise. On july 8, it paid the full amount due. The amount of the cash paid on july 8 equals: a) 330 b) 2610 c)2617 d) 2670
Answer #1

terms 2/10 net 30 means that the buyer is eligible for 1 2% discount if the payment is made in 10 days else the entire payment has to be made within 30 days

Net purchase = 3000 - 330 = 2670

Discount since paid within 3 days of purchase = 2670 *2% = 53.40

Net paid = 2670 - 53.40 = 2617